Requirements: Please finish the following questions using Excel with appropriate formulas; Please clearly specify your final answers (i.e. indicate the final decision given the criteria); All work should be completed in one single Excel file. Please refer to the last page for how to insert separate worksheets in Excel; Submission must be uploaded through Blackboard. When submit, please name your file with the format: A1_First name_Last name.

Q2.The XYZ Company is considering adding a new service facility among several possible alternatives including a child care center, a swimming pool, new locker rooms and showers, a health-oriented food court, and a spa. The following payoff table summarizes the potential returns (based on costs and increased enrollments) for each alternative service facility given three future levels of demand.

Demand

Service Facility

Poor

Moderate

High

Child care center

$16,000

$25,000

$43,000

Swimming pool

-$75,000

$27,000

$71,000

New lockers & showers

$11,500

$37,000

$56,000

Food court

-$31,000

$17,000

$87,000

Spa

$6,000

$25,600

$33,500

Determine the best decision using the following criteria.

  1. Maximax
  2. Maximin
  3. Hurwicz (α= 0.56)
  4. Equal likelihood
  5. Expected value (Poor: 0.18; Moderate: 0.56; High: 0.26)

==========================================================================

  • How to insert separate worksheets in Excel
Answer #1

The below screenshot represent the consolidated solution for the given payoff table:

EMV Row Min Row Max Probabilities Child care center Swimming Pool New Lockers & Showers Food Court Spa Poor 0.18 16000 -75000

A) Maximax Criterion:

The Column Row Max shows the maximum value of each of the five decision alternatives considering all the three states of nature.

Maximax = Max(Row Column elements) = Max{D1, D2,D3,D4,D5} = Max{43000,71000,56000,87000,33500}

= 87000 (Under Maximax, We choose Foodcourt)

B) Maximin Criterion:

The column RowMin specifies the minimum value of each of the alternatives and it is the conservative approach.

Maximin = Max(Row Min) = Max{16000, -75000, 11500, -31000, 6000} = 16000

Under Maximin, We choose Child Care center.

C) Hurwicz (At alpha = 0.56)

Hurwiz value = alpha * Maximum + (1-alpha)* Minimum    (Max & Min values For each decision alternative)

child care center = 0.56*43000 + (1-0.56)*16000 = 31120

swimming pool = 0.56*71000 + (1-0.56)*-75000 = 6760

New Lockers = 0.56*56000 + (1-0.56)*11500 = 36420

Food court = 0.56*87000 + (1-0.56)*-31000 = 35080

Spa = 0.56*33500 + (1-0.56)*6000 = 21400

The highest return as per Hurwicz is for New lockers & showers with the return of $36,420

D) Expected Monetary value = Sumproduct (Probability of state, Return)

EMV for child care center = 0.18*16000 + 0.56*25000 + 0.26*43000 = 28060

EMV for swimming pool = 0.18*-75000 + 0.56*27000 + 0.26*71000 = 20080

Similarly,

EMV for New Lockers = 37350

EMV for Food court = 26560

EMV for Spa = 24126

The largest of the EMV values gives the expected return for the best alternative. It is for New lockers & showers at $37350